OpenStudy (anonymous):

What is the approximate volume of the cone? Use 3.14 for π. A. 4239 cm3 B. 6359 cm3 C. 1272 cm3 D. 2120 cm3

pooja195 (pooja195):

\[\pi~r^2 \frac{ h }{ 3 }\]

pooja195 (pooja195):

plug in the numbers you will have your answer :)

OpenStudy (anonymous):

15 times 15 = 225 times 9 = 2025 and 2025 divide by 3 = 675

OpenStudy (anonymous):

am i doing this right? @pooja195

OpenStudy (anonymous):

so D

pooja195 (pooja195):

yep :)

OpenStudy (anonymous):

Can you show me how to do it

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

\[(3.14)\times15^2~~\frac{ 9 }{ 3 }=2119.5\rightarrow~2120~(rounded)\]